# Online Detecting LLM-Generated Texts via Sequential Hypothesis Testing by Betting

### v1

@@ Appendix E Proof of Proposition 1
By definition of the rejection time and that $\{\tau=\infty\}\subseteq\{\tau\geq t\}$ for all $t\geq 1$, we know $\{\tau>t\}\subseteq\{W_{t}<\frac{1}{\alpha}\}$ and $P(\tau=\infty)\leq\liminf_{t\to\infty}P(\tau>t)\leq\liminf_{t\to\infty}P(W_{t}<1/\alpha)$. By the second inequality of (42),
@@
It is almost surely that $S_{t}/t=\frac{1}{t}\sum_{i=1}^{t}(\phi({x}_{t})-\phi({y}_{t}))$ converges to $(\mu_{x}-\mu_{y})$ as $t\rightarrow\infty$, according to the Strong Law of Large Numbers. We recall that under $H_{1}:\mu_{x}-\mu_{y}\neq 0$. On the other hand, $\frac{4d_{*}^{4}+4d_{*}^{3}}{2d_{*}-1}\cdot\frac{\log(t^{\frac{1}{2\gamma}}d_{*}^{\frac{1}{\gamma}}/\alpha)}{t}\to 0$ as $t\rightarrow\infty$. Thus, if we let $a_{t}$ be the event that $\exp\left(\frac{2d_{*}-1}{4d_{*}^{2}}\cdot\frac{S_{t}^{2}}{td_{*}^{2}+td_{*}}\right)<(td_{*}^{2})^{\frac{1}{2\gamma}}/\alpha$ , we see that $\bm{1}_{a_{t}}\to 0$ almost surely. By the dominated convergence theorem,
@@
$\Rightarrow E_{t}$$\subseteq\left\{{S_{t}^{2}}<\frac{4d_{*}^{2}}{2d_{*}-1}\left(Q_{t}+|S_{t}|\right)\left(\log(1/\alpha)-\log(1/Q_{t}^{\frac{1}{2\gamma}})\right)\right\}\quad\left(\text{since }\left\lvert S_{t}\right\rvert=\left\lvert\sum_{i=1}^{t}g_{i}\right\rvert\leq\sum_{i=1}^{t}\left\lvert g_{i}\right\rvert\right)$
$\subseteq\left\{{S_{t}^{2}}<\frac{4d_{*}^{2}}{2d_{*}-1}\left(Q_{t}+\sum_{i=1}^{t}|g_{i}|\right)\left(\log(1/\alpha)-\log(1/Q_{t}^{\frac{1}{2\gamma}})\right)\right\}.$ (45)
@@ Appendix F Proof of Proposition 2
Algorithm 3 Online Detecting LLMs via Online Optimization and Betting for the Composite Hypothesis Testing
@@ Appendix G Experiment Results of Detecting 2024 Olympic News or Machine-generated News
(f) Results for detecting the source of text $y_{t}$, which is 2024 Olympic news or news generated by PaLM 2, with human-written text $x_{t}$ sampled from XSum. The scoring model used is Gemma-2B.
@@
(f) Results for detecting the source of text $y_{t}$, which is 2024 Olympic news or news generated by PaLM 2, with human-written text $x_{t}$ sampled from XSum. The scoring model used is Gemma-2B.
@@
Table 1: Values of $\Delta$, derived by applying different scoring models and various score functions to text $x_{t}$ from Xsum and $y_{t}$ from 2024 Olympic news or machine-generated news. The value of $\Delta$ is calculated according to $\Delta=\left\lvert\left({\sum_{i=1}^{500}\phi(x_{i})}\right)/{500}-\left({\sum_{j=1}^{500}\phi(y_{j})}\right)/{500}\right\rvert$, where $x_{i}$ is score of the $i$-th text from XSum, $y_{j}$ is score of the $j$-th text from the source to be detected. Every two columns starting from the third column represent the $\Delta$ values under $H_{1}$ and $H_{0}$ for each test scenario. For instance, in calculating $\Delta$ for the third column, $y_{j}$ represents the $j$-th fake news generated by Gemini-1.5-Flash based on pre-tokens of Olympic 2024 news. For the fourth column, $y_{j}$ refers to the $j$-th 2024 Olympic news articles. Values in Column “Human, Human" are also used to set $\epsilon$ values for tests in Scenario 1.
@@
Table 2: Values of $d_{t}$ used in Scenario 1 with the assumption that we know the range of $g_{t}=\lvert\phi(x_{t})-\phi(y_{t})\rvert$ beforehand. Every two columns starting from the third column represent the $d_{t}$ values for each $t$ used under $H_{1}$ and $H_{0}$ in each test scenario. For instance, the values in the third column are calculated as $\max_{i,j\leq 500}\lvert\phi(x_{i})-\phi(y_{j})\rvert$, where $\phi(x_{i})$ is the score of $i$-th XSum text and $\phi(y_{j})$ is the score of the $j$-th text generated by Gemini-1.5-Flash, prompted by pre-tokens of 2024 Olympic news. This calculation ensures that $d_{t}\geq\lvert\phi(x_{t})-\phi(y_{t})\rvert$ for any time point $1\leq t\leq 500$. Similarly, the fourth column calculates the maximum difference between the scores of all XSum texts and texts sample of 2024 Olympic news.. The derived $d_{t}$ values is then used to define the domain of $\theta_{t}$ in our algorithm, where $\theta_{t}\in[-1/2d_{t},1/2d_{t}]$.
@@ Appendix H Experiment Results of detecting texts from three domains
(a) Results when $x_{t}$ is sampled from XSum and $y_{t}$ is sampled from XSum or generated by GPT-3 based on XSum.
(b) Results when $x_{t}$ is sampled from XSum and $y_{t}$ is sampled from XSum or generated by ChatGPT based on XSum.
(c) Results when $x_{t}$ is sampled from XSum and $y_{t}$ is sampled from XSum or generated by GPT-4 based on XSum.
(d) Results when $x_{t}$ is sampled from Writing and $y_{t}$ is sampled from Writing or generated by GPT-3 based on Writing.
(e) Results when $x_{t}$ is sampled from Writing and $y_{t}$ is sampled from Writing or generated by ChatGPT based on Writing.
(f) Results when $x_{t}$ is sampled from Writing and $y_{t}$ is sampled from Writing or generated by GPT-4 based on Writing.
(g) Results when $x_{t}$ is sampled from PubMed and $y_{t}$ is sampled from PubMed or generated by GPT-3 based on PubMed.
(h) Results when $x_{t}$ is sampled from PubMed and $y_{t}$ is sampled from PubMed or generated by ChatGPT based on PubMed.
(i) Results when $x_{t}$ is sampled from PubMed and $y_{t}$ is sampled from PubMed or generated by GPT-4 based on PubMed.
@@
Figure 9: Test results: mean rejection times (under $H_{1}$) and FPRs (under $H_{0}$) under each significance level $\alpha$ using 10 score functions, with texts $x_{t}$ and $y_{t}$ from the same domain. There are 3 source models: GPT-3, ChatGPT and GPT-4.
(a) Results for $x_{t}$ sampled from XSum and $y_{t}$ sampled from Writing or generated by GPT-3 based on Writing.
(b) Results for $x_{t}$ sampled from XSum and $y_{t}$ sampled from Writing or generated by ChatGPT based on Writing.
(c) Results for $x_{t}$ sampled from XSum and $y_{t}$ sampled from Writing or generated by GPT-4 based on Writing.
(d) Results for $x_{t}$ sampled from XSum and $y_{t}$ sampled from PubMed or generated by GPT-3 based on PubMed.
(e) Results for $x_{t}$ sampled from XSum and $y_{t}$ sampled from PubMed or generated by ChatGPT based on PubMed.
(f) Results for $x_{t}$ sampled from XSum and $y_{t}$ sampled from PubMed or generated by GPT-4 based on PubMed.
(g) Results for $x_{t}$ sampled from Writing and $y_{t}$ sampled from PubMed or generated by GPT-3 based on PubMed.
(h) Results for $x_{t}$ sampled from Writing and $y_{t}$ sampled from PubMed or generated by ChatGPT based on PubMed.
(i) Results for $x_{t}$ sampled from Writing and $y_{t}$ sampled from PubMed or generated by GPT-4 based on PubMed.
@@
Figure 10: Test results: mean rejection times (under $H_{1}$) and FPRs (under $H_{0}$) under each significance level $\alpha$ using 10 score functions, with texts $x_{t}$ and $y_{t}$ from different domains. There are 3 source models: GPT-3, ChatGPT and GPT-4.

### v2

@@ Appendix E Proof of Proposition 1
By definition of the rejection time, we know $\{\tau>t\}\subseteq\{W_{t}<\frac{1}{\alpha}\}$. Because $\{\tau=\infty\}\subseteq\{\tau\geq t\}$ for all $t\geq 1$, we can ultimately get $P(\tau=\infty)\leq\liminf_{t\to\infty}P(\tau>t)\leq\liminf_{t\to\infty}P(W_{t}<1/\alpha)$. By the second inequality of (40),
@@
It is almost surely that $S_{t}/t=\frac{1}{t}\sum_{i=1}^{t}(\phi({x}_{t})-\phi({y}_{t}))$ converges to $(\mu_{x}-\mu_{y})$ as $t\rightarrow\infty$, according to the Strong Law of Large Numbers. We recall that under $H_{1}:\mu_{x}-\mu_{y}\neq 0$. On the other hand, $\frac{4d_{*}^{4}+4d_{*}^{3}}{2d_{*}-1}\cdot\frac{\log(t^{\frac{1}{2\gamma}}d_{*}^{\frac{1}{\gamma}}/\alpha)}{t}\to 0$ as $t\rightarrow\infty$. Thus, if we let $a_{t}$ be the event that $\exp\left(\frac{2d_{*}-1}{4d_{*}^{2}}\cdot\frac{S_{t}^{2}}{td_{*}^{2}+td_{*}}\right)<(td_{*}^{2})^{\frac{1}{2\gamma}}/\alpha$ , we see that $\mathcal{1}_{a_{t}}\to 0$ almost surely. By the dominated convergence theorem,
@@
$\Rightarrow E_{t}$$\subseteq\left\{{S_{t}^{2}}<\frac{4d_{*}^{2}}{2d_{*}-1}\left(Q_{t}+|S_{t}|\right)\left(\log(1/\alpha)-\log(1/Q_{t}^{\frac{1}{2\gamma}})\right)\right\}$
$\subseteq\left\{{S_{t}^{2}}<\frac{4d_{*}^{2}}{2d_{*}-1}\left(Q_{t}+\sum_{i=1}^{t}|g_{i}|\right)\left(\log(1/\alpha)-\log(1/Q_{t}^{\frac{1}{2\gamma}})\right)\right\}\quad\left(\text{Because }\left\lvert S_{t}\right\rvert=\left\lvert\sum_{i=1}^{t}g_{i}\right\rvert\leq\sum_{i=1}^{t}\left\lvert g_{i}\right\rvert\right).$ (43)
@@ Appendix F Proof of Proposition 2
Algorithm 3 Online Detecting LLMs via Online Optimization and Betting for the Composite Hypotheses Testing
@@ Appendix G Experiment Results of Detecting 2024 Olympic News or Machine-generated News
Our generation process for fake news is guided by Mitchell et al. [2023], Bao et al. [2023], Su et al. [2023]. Specifically, we use the T5 tokenizer to process each human-written news article to retrieve the first 30 tokens as {prefix}. Then, we initiate the generation process by sending the following messages to the model service, such as: "You are a News writer. Please write an article with about 150 words starting exactly with {prefix}."
@@
(f) Results for detecting the source of text $y_{t}$, which is 2024 Olympic news or news generated by PaLM 2, with text $x_{t}$ sampled from XSum. The scoring model used is Gemma-2B.
@@
(f) Results for detecting the source of text $y_{t}$, which is 2024 Olympic news or news generated by PaLM 2, with text $x_{t}$ sampled from XSum. The scoring model used is Gemma-2B.
@@
Table 1: Values of $\Delta$, which are calculated according to $\Delta=\left\lvert\left({\sum_{i=1}^{500}\phi(x_{i})}\right)/{500}-\left({\sum_{j=1}^{500}\phi(y_{j})}\right)/{500}\right\rvert$, where $x_{i}$ is score of the $i$-th text from XSum, $y_{j}$ is score of the $j$-th text from the source to be detected. Every two columns starting from the third column represent the $\Delta$ values under $H_{1}$ and $H_{0}$ for each test scenario. For instance, in calculating $\Delta$ for the third column, $y_{j}$ represents the $j$-th fake news generated by Gemini-1.5-Flash based on pre-tokens of Olympic 2024 news. For the fourth column, $y_{j}$ refers to the $j$-th 2024 Olympic news articles. Values in Column “Human, Human" are also used to set $\epsilon$ values for tests in Scenario 1.
@@
Table 2: Values of $d_{t}$ used in Scenario 1, for which we assume that the range of $g_{t}=\lvert\phi(x_{t})-\phi(y_{t})\rvert$ is known beforehand. Specifically, $d_{t}$ is calculated as $\max_{i,j\leq 500}\lvert\phi(x_{i})-\phi(y_{j})\rvert$, where $\phi(x_{i})$ is the score of $i$-th XSum text and $\phi(y_{j})$ is the score of the $j$-th text generated by Gemini-1.5-Flash, prompted by pre-tokens of 2024 Olympic news. This calculation ensures that $d_{t}\geq\lvert\phi(x_{t})-\phi(y_{t})\rvert$ for any time point $1\leq t\leq 500$. Every two columns starting from the third column represent the $d_{t}$ values for each $t$ used under $H_{1}$ and $H_{0}$ in each test scenario. For instance, the values in the third column Similarly, the fourth column calculates the maximum difference between the scores of all XSum texts and texts sample of 2024 Olympic news.. The derived $d_{t}$ values is then used to define the domain of $\theta_{t}$ in our algorithm, where $\theta_{t}\in[-1/2d_{t},1/2d_{t}]$.
@@ Appendix H Experiment Results of detecting texts from three domains
(a) Same domain: $x_{t}$ is sampled from XSum, and $y_{t}$ is sampled from XSum or generated by GPT-3 based on XSum.
(b) Same domain: $x_{t}$ is sampled from XSum, and $y_{t}$ is sampled from XSum or generated by ChatGPT based on XSum.
(c) Same domain: $x_{t}$ is sampled from XSum, and $y_{t}$ is sampled from XSum or generated by GPT-4 based on XSum.
(d) Same domain: $x_{t}$ is sampled from Writing, and $y_{t}$ is sampled from Writing or generated by GPT-3 based on Writing.
(e) Same domain: $x_{t}$ is sampled from Writing, and $y_{t}$ is sampled from Writing or generated by ChatGPT based on Writing.
(f) Same domain: $x_{t}$ is sampled from Writing, and $y_{t}$ is sampled from Writing or generated by GPT-4 based on Writing.
(g) Same domain: $x_{t}$ is sampled from PubMed, and $y_{t}$ is sampled from PubMed or generated by GPT-3 based on PubMed.
(h) Same domain: $x_{t}$ is sampled from PubMed, and $y_{t}$ is sampled from PubMed or generated by ChatGPT based on PubMed.
(i) Same domain: $x_{t}$ is sampled from PubMed, and $y_{t}$ is sampled from PubMed or generated by GPT-4 based on PubMed.
@@
Figure 9: Test results: mean rejection times (under $H_{1}$) and FPRs (under $H_{0}$) of each significance level $\alpha$ for 10 score functions, with texts $x_{t}$ and $y_{t}$ from the same domain. There are 3 source models: GPT-3, ChatGPT and GPT-4.
(a) Different domains: $x_{t}$ is sampled from XSum, and $y_{t}$ is sampled from Writing or generated by GPT-3 based on Writing.
(b) Different domains: $x_{t}$ is sampled from XSum, and $y_{t}$ is sampled from Writing or generated by ChatGPT based on Writing.
(c) Different domains: $x_{t}$ is sampled from XSum, and $y_{t}$ is sampled from Writing or generated by GPT-4 based on Writing.
(d) Different domains: $x_{t}$ is sampled from XSum, and $y_{t}$ is sampled from PubMed or generated by GPT-3 based on PubMed.
(e) Different domains: $x_{t}$ is sampled from XSum, and $y_{t}$ is sampled from PubMed or generated by ChatGPT based on PubMed.
(f) Different domains: $x_{t}$ is sampled from XSum, and $y_{t}$ is sampled from PubMed or generated by GPT-4 based on PubMed.
(g) Different domains: $x_{t}$ is sampled from Writing, and $y_{t}$ is sampled from PubMed or generated by GPT-3 based on PubMed.
(h) Different domains: $x_{t}$ is sampled from Writing, and $y_{t}$ is sampled from PubMed or generated by ChatGPT based on PubMed.
(i) Different domains: $x_{t}$ is sampled from Writing, and $y_{t}$ is sampled from PubMed or generated by GPT-4 based on PubMed.
@@
Figure 10: Test results: mean rejection times (under $H_{1}$) and FPRs (under $H_{0}$) of each significance level $\alpha$ for 10 score functions, with texts $x_{t}$ and $y_{t}$ from different domains. There are 3 source models: GPT-3, ChatGPT and GPT-4.
@@
Appendix I Addressing Parameter Estimation Challenges and Mixed LLMs or Sources Tasks
Estimate Parameters Based on More Samples. Better parameter estimation can enhance the performance of our algorithm. For instance, in the previous experiments, we used the first $10$ samples to estimate parameters $d_{t}$ and $\epsilon$. Here, we give an example to show that using a larger sample size for estimation could possibly yield better results. Specifically, using $20$ samples for estimation, with test begins at the $21$-th time step, could lead to improved algorithm performance.
Figure 11 demonstrates that when parameters are estimated with more samples from the initial time steps, all score functions maintain False Positive Rates (FPRs) below the specified significance levels. Additionally, almost all functions identify the LLM source more quickly compared to when fewer samples are used for estimation. This example indicates the potential benefits of using more extensive data for parameter estimation in enhancing the effectiveness of our approach.
(a) Averaged test results for estimating parameters based on first $10$ samples.
(b) Averaged test results for estimating parameters based on first $20$ samples.
Figure 11: Comparison of different durations in the initial stage for parameter estimation in Scenario 2. Here, text $x_{t}$ is sampled from XSum and $y_{t}$ is from 2024 Olympic news or machine-generated news, across three source models. The scoring model is Neo-2.7. Better parameter estimation can enhance the performance of our algorithm. In the previous experiments, we used the first $10$ samples to estimate parameters $d_{t}$ and $\epsilon$, as shown in (a). The subfigure (b) suggests that a longer duration of for parameter estimation could possibly yield better results, where we emphasize that the test begins at $t=21$ after estimating the parameters.
Tasks for a Mixture of LLMs or Sources. Our method can be extended to additional tasks. Its fundamental goal in sequential hypothesis testing is to determine whether texts from an unknown source originate from the same distribution as those in a prepared human text dataset, where we consider mean value as the statistical metric.
Even if texts from the LLM source are produced by various LLMs, they still satisfy the alternative hypothesis, which means that our statistical guarantees remain valid and the algorithm could continue to perform effectively. The results are illustrated in Figure 12(a).
When the unknown source publishes both human-written and LLM-generated texts, our method can effectively address this scenario. Here, the null hypothesis assumes that all texts from the unknown source are human-written. In contrast, the alternative hypothesis posits that not all texts are human-written, which indicates the presence of texts generated by LLMs. Figure 12(b) demonstrates that our algorithm, equipped with nearly all score functions, consistently performs well in this new context.
(a) Task 1: When the LLM source posts texts generated by different LLMs (under $H_{1}$). Specifically, the sequence consists of $100$ texts generated by Gemini-1.5-Pro, $200$ texts generated by Gemini-1.5-Flash, and $200$ texts generated by PaLM 2.
(b) Task 2: When the unknown source posts a mixture of human-written texts and LLM-generated texts (under $H_{1}$). Specifically, the sequence consists of $200$ texts written by human, and $300$ texts generated by PaLM 2.
Figure 12: (Extension to other settings) (a) Results when the sequence of texts $y_{t}$ are produced by various LLMs instead of a single one. (b) Results under the setting that the null hypothesis corresponds to the case that all the texts from the unknown source are human-written, while the alternative hypothesis corresponds to the one that not all $y_{t}$ are human-written.
The above results reflect a real-world scenario where parameter values are estimated from the first $10$ samples. The performance of our method could be further enhanced with prior knowledge of the parameters.